directions: convert the following units to the indicated units
1) 20 liters/ hour to milliliters/minute.
2) 80 mAs to Amp Hours

Answer by stanbon(75887)

convert the following units to the indicated units
1) 20 liters/ hour to milliliters/minute.
---
(20 L/hr)(1000mL/L)(60 min/hr) = [(20*1000*60)]/[1*1*1] (mL/minute)
= 1,200,000 (mL/minute)
================================
2) 80 mAs to Amp Hours
(80 mA sec)(Amp/1000 mA)(hr/3600 sec)
= [80*1*1]/[1*1000*3600)(Amp hrs)
= 80/3600000 Amp hrs
= 0.00002222.. Amp hrs
============================
Cheers,
Stan H.